OpenStudy (anonymous):

Find the next three terms of the sequence. –2, –12, –72, –432, … • –1,728, –6,912, –27,648 • –1,728, –10,368, –62,208 • –2,592, –10,368, –41,472 • –2,592, –15,552, –93,312 My answer is B check my work plz

OpenStudy (welshfella):

to get the next term multiply by 6 -2*6 = -12 -12*6 = -72 etc

OpenStudy (anonymous):

D

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Thanks!

OpenStudy (welshfella):

yw